Question three: A 165.5 L kiln is used for vitrifying ceramics. It is currently operating at 1091.0 ∘C , and the pressure is 1.075 atm . How many moles of air molecules are within the confines of the kiln?

Question four:

A 24.20 L gas cylinder has been filled with 3.922 moles of gas. You measure the pressure to be 3.330 atm . What is the temperature inside the tank?

Express your answer numerically in degrees Celsius. [Pay attention to the number of decimal places you are allowed in your answer!]
Answer #1

V = 165.5 L

T = 1091C or 1364C

P = 1.075 atm

n = ?

find moles

apply idel gas

PV = nRT

n = PV/(RT) = 1.074*165.5/0.082/1364 = 1.589 mol of air

Q4

V = 24.2 L

n = 3.922 mol

P = 3.33 atm

T = ?

PV = nRT

T = PV/(nR) = 3.33*24.20/(3.922*0.082) = 250.6 K

T = 250.6K - 273 = -22.4 C